A multimode step index fiber has a material dispersion parameter of 250 ps 

nm−1km−l at 1.2 μm. Estimate the pulse broadening for a 10 kilometer long fiber 

with an LED source of spectral width 50 nm.


1
Expert's answer
2021-07-07T08:39:16-0400

s

elative refractive index ∆ = 2 % = 0.02

Core refractive index n1 = 1.48

Link length L = 5 km = 5000 m

Maximum optical bandwidth "BW = 3 MHz = 3\times 10^6 \;Hz"

Material dispersion parameter M = 80 ps nm-1 km-1

Spectral width due to waveguide dispersion "\u03c3_\u03bb = 1"

The expression of the rms pulse broadening due to chromatic dispersion τ:

"τ ≈ \frac{L(NA)^2}{2n_{1}c}"

NA is the numerical aperture, n1 is core refractive index

is the velocity of light "3\times 10^8\; m/s"

The expression for numerical aperture NA:

"NA= n_1\sqrt{2∆}"

∆ is the relative refractive index

"τ ≈ \frac{L(n_1\sqrt{2∆})^2}{2n_1c}"

"τ ≈ \frac{5000(1.48\sqrt{2\times0.02})^2}{2\times1.48\times 3\times10^8}"

"τ ≈ 0.493 \;μs"

The expression to calculate the rms pulse broadening per km

"\frac{τ}{L} = \frac{0.493\; μs}{5\; km} = 98.6\; \frac{ns}{km}"

The expression to calculate spectral width of source:

"σ_m = σ_λ\times M\times L"

"σ_m = 1\times 80\times 5 = 0.4 \; \frac{ns}{km}"